OpenStudy (anonymous):

Find the LCM of y^3+10y^2+25y and y^2-10y.

jimthompson5910 (jim_thompson5910):

y^3+10y^2+25y = y(y^2+10y+25) = y(y+5)(y+5) y^2-10y = y(y-10) Now collect the unique factors and multiply LCM = y(y+5)(y+5)(y-10) Note: you collect the factors that occur the most as well

OpenStudy (anonymous):

I'm beginning to understand this a bit more. Thank you for your help. I have another one that looks a bit different to me. It is, "Find the LCM of 14x^2+42x and 7x^2+49x+84.

jimthompson5910 (jim_thompson5910):

14x^2+42x = 14x(x+3) = 2*7x(x+3) 7x^2+49x+84 = 7(x^2+7x+12) = 7(x+3)(x+4) Now collect unique and most occurring factors to get LCM = 2*7*x*(x+3)(x+4) = 14x(x+3)(x+4) So the LCM is 14x(x+3)(x+4)

OpenStudy (anonymous):

You've been a great help. Thank you!